OpenStudy (anonymous):

equation for a line passing through the point (5,1) and (-2,3)

OpenStudy (haseeb96):

first find the slope from these points m=y2-y1 / x2 -x1 let (x1,y1) =(5,1) (x2,y2)=(-2,3) then m= 3-1 / -2-5 m=2/ -7 m=- 2/7 now u have slope and points so take any one point from the given data and put both slope and the point into this formula( point slope formula) (x1,y1)=(5,1) (y-y1)=m(x-x1) y-1= - 2 /7 ( x- 5) 7y -7 =-2x+10 7y +2x -7 -10=0 2x+7y -17=0 is the equation of the line have u understood it? if u have any question then ask me ? i will try my best to help u